\begin{matrix}c \\ ◯ \end{matrix} R.J. van Glabbeek, C. Author & Y.S. Else This work is licensed under the Creative Commons Attribution License.

An Example of a Paper with a Rather Large Title-to-Content Ratio

Rob van Glabbeek

NICTA Sydney, Australia School of Computer Science and Engineering University of New South Wales^{∗} Sydney, Australia

                 rvg@cs.stanford.edu

Co Author Yet S. Else

Stanford Univeristy California, USA

                                                    is@gmail.com       somebody@else.org

This is a sentence in the abstract. This is another sentence in the abstract. This is yet another sentence in the abstract. This is the final sentence in the abstract.

1 Introduction

The optional arguments of \documentclass{eptcs} are

   • at most one of adraft, submission or preliminary,
   • at most one of publicdomain or copyright,
   • and optionally creativecommons,

– possibly augmented with

           ∗ noderivs
           ∗ or sharealike,

– and possibly augmented with noncommercial.

We use adraft rather than draft so as not to confuse hyperref. The style-file option submission is for papers that are submitted to \event, where the value of the latter is to be filled in in line 2 of the tex-file. Use preliminary only for papers that are accepted but not yet published. The final version of your paper that is to be uploaded at the EPTCS website should have none of these style-file options.

By means of the style-file option creativecommons authors equip their paper with a Creative Commons license that allows everyone to copy, distribute, display, and perform their copyrighted work and derivative works based upon it, but only if they give credit the way you request. By invoking the additional style-file option noderivs you let others copy, distribute, display, and perform only verbatim copies of your work, but not derivative works based upon it. Alternatively, the sharealike option allows others to distribute derivative works only under a license identical to the license that governs your work. Finally, you can invoke the option noncommercial that let others copy, distribute, display, and perform your work and derivative works based upon it for noncommercial purposes only.

Authors’ (multiple) affiliations and emails use the commands \institute and \email. Both are optional. Authors should moreover supply \titlerunning and \authorrunning, and in case the copyrightholders are not the authors also \copyrightholders. As illustrated above, heuristic solutions may be called for to share affiliations. Authors may apply their own creativity here [1].

Submitted to:

SOS 2007

∗A fine university.

   EPTCS recommends using \documentclass[copyright,creativecommons]{eptcs}.

Additionally, the title should be set in title case, meaning that major words start with a capital letter, and only articles, prepositions and conjunctions appear in lower case.

Exactly 46 lines fit on a page. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 6 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 11 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details.

Here starts a new paragraph. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 16 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 21 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 26 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 31 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 36 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 41 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. 46 The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details. The rest is like any normal L^{A}TEX article. We will spare you the details.

2 Ancillary files

Authors may upload ancillary files to be linked alongside their paper. These can for instance contain raw data for tables and plots in the article, or program code. Ancillary files are included with an EPTCS submission by placing them in a directory anc next to the main latex file. See also https://arxiv. org/help/ancillary_files. Please add a file README in the directory anc, explaining the nature

of the ancillary files, as in http://eptcs.org/paper.cgi?226.21.

3 Prefaces

Volume editors may create prefaces using this very template, with \title{Preface} and \author{}.

4 Bibliography

We request that you use \bibliographystyle{eptcs} [2], or one of its variants eptcsalpha, eptcsini or eptcsalphaini [2]. Compared to the original L^{A}TEX \bibliographystyle{plain}, it ignores the field month, and uses the extra bibtex fields eid, doi, eprint and url. The first is for electronic identifiers (typically the number n indicating the n^{th} paper in an issue) of papers in electronic journals that do not use page numbers. The other three are to refer, with life links, to electronic incarnations of the paper.

DOIs Almost all publishers use digital object identifiers (DOIs) as a persistent way to locate electronic publications. Prefixing the DOI of any paper with http://dx.doi.org/ yields a URI that resolves to the current location (URL) of the response page^{1} of that paper. When the location of the response page changes (for instance through a merge of publishers), the DOI of the paper remains the same and (through an update by the publisher) the corresponding URI will then resolve to the new location. For that reason a reference ought to contain the DOI of a paper, with a life link to the corresponding URI, rather than a direct reference or link to the current URL of publisher’s response page. This is the rôle of the bibtex field doi. EPTCS requires the inclusion of a DOI in each cited paper, when available.

1Nowadays, papers that are published electronically tend to have a response page that lists the title, authors and abstract of the paper, and links to the actual manifestations of the paper (e.g. as dvi- or pdf-file). Sometimes publishers charge money to access the paper itself, but the response page is always freely available.

DOIs of papers can often be found through http://www.crossref.org/guestquery;^{2} the second method Search on article title, only using the surname of the first-listed author, works best. Other places to find DOIs are DBLP and the response pages for cited papers (maintained by their publishers).

The bibtex fields eprint and url Often an official publication is only available against payment, but as a courtesy to readers that do not wish to pay, the authors also make the paper available free of charge at a repository such as arXiv.org. In such a case it is recommended to also refer and link to the URL of the response page of the paper in such a repository. This can be done using the bibtex fields eprint or url. The latter field should not be used to duplicate information that is also provided through doi or eprint. You can find archival-quality URL’s for most recently published papers in DBLP, but please suppress repetition of DOI or eprint information though url. In fact, it is often useful to check your references against DBLP records anyway, or just find them there in the first place.

Typesetting DOIs and URLs When using L^{A}TEX rather than pdflatex to typeset your paper, by default no linebreaking within long URLs is allowed. This leads often to very ugly output, that moreover is different from the output generated when using pdflatex. This problem is repaired when invoking \usepackage{breakurl}: it allows linebreaking within links and yield the same output as obtained by default with pdflatex. When invoking pdflatex, the package breakurl is ignored.

The package \usepackage{underscore} is recommended to deal with underscores in DOIs. This is not needed when using \usepackage{breakurl} and not pdflatex.

References to papers in the same EPTCS volume To refer to another paper in the same volume as your own contribution, use a bibtex entry with

                           series = {\thisvolume{5}},

where 5 is the submission number of the paper you want to cite. You may need to contact the author, volume editors or EPTCS staff to find that submission number; it becomes known (and unchangeable) as soon as the cited paper is first uploaded at EPTCS. Furthermore, omit the fields publisher and volume. Then in your main paper you put something like:

\providecommand{\thisvolume}[1]{this volume of EPTCS, Open Publishing Association}

This acts as a placeholder macro-expansion until EPTCS software adds something like

\newcommand{\thisvolume}[1]{{\eptcs} 157\opa, pp 45--56, doi:...},

where the relevant numbers are pulled out of the database at publication time. Here the newcommand wins from the providecommand, and \eptcs resp. \opa expand to

\sl Electronic Proceedings in Theoretical Computer Science  and
, Open Publishing Association  .

Hence putting \def\opa{} in your paper suppresses the addition of a publisher upon expansion of the citation by EPTCS. An optional argument like

                        series = {\thisvolume[EPTCS]{5}},

overwrites the value of \eptcs.

2For papers that will appear in EPTCS and use \bibliographystyle{eptcs} there is no need to find DOIs on this website, as EPTCS will look them up for you automatically upon submission of a first version of your paper; these DOIs can then be incorporated in the final version, together with the remaining DOIs that need to found at DBLP or publisher’s webpages.

R.J. van Glabbeek, C. Author & Y.S. Else 5

References

[1] Stack Exchange: gernot answering effeffe (2016): Centering issues with multiple authors with the same

affiliation, EPTCS format. Available at http://tex.stackexchange.com/q/344794/63917.

[2] R.J. van Glabbeek (2010): The EPTCS bibliography style. Available at http://biblio.eptcs.org.

[3] E.D.I. Thor & E. di Thor, editors (1987): Title B1: A Book with editors, second edition. Series 21, Publisher, Address, doi:10.4204/EPTCS. arXiv:2042.12345. Note.

[4] E.D.I. Thor & E. di Thor, editors (1987): Title B2: An Inbook with editors, second edition, type (of chapter) II, pp. 1–999. Series 443, Publisher, Address, doi:10.4204/EPTCS. arXiv:2042.12345. Note.

[5] E.D.I. Thor & E. di Thor, editors (1987): Title B3: A Proceedings. Series 443, Organization, Publisher,

Address, doi:10.4204/EPTCS. Available at https://ntrs.nasa.gov/citations/20205011579. Note.

[6] A. Ut, H. Or & Co Author (1987): Title A: An Article. Jounal 443(21):4, doi:10.4204/EPTCS. arXiv:2042. 12345. Note.

[7] A. Ut, H. Or & Co Author (1987): Title B1: A Book with authors, second edition. Series 443, Publisher,

Address, doi:10.4204/EPTCS. Available at https://ntrs.nasa.gov/. Note.

[8] A. Ut, H. Or & Co Author (1987): Title B2: An Inbook with authors, second edition, type (of chapter) II, pp. 1–999. Series 443, Publisher, Address, doi:10.4204/EPTCS. Available at https://ntrs.nasa.gov/

citations/20205011579. Note.

[9] A. Ut, H. Or & Co Author (1987): Title C1: An Inproceedings. In E.D.I. Thor & E. di Thor, editors: Booktitle, Series 443, Organization, Publisher, Address, pp. 1–999, doi:10.4204/EPTCS. arXiv:2042.12345. Note.

[10] A. Ut, H. Or & Co Author (1987): Title C2: An Incollection. In E.D.I. Thor & E. di Thor, editors: Booktitle, second edition, type II, Series 443, Publisher, Address, pp. 1–999, doi:10.4204/EPTCS. Available at

https://ntrs.nasa.gov/citations/20205011579. Note.

[11] A. Ut, H. Or & Co Author (1987): Title R: A Techreport. Type (of report) 21, Institution, Address, doi:10.

4204/EPTCS. arXiv:2042.12345. Note.

[12] A. Ut, H. Or & Co Author (1987): Title R1: A Techreport. Technical Report 21, Institution, Address, doi:10.

4204/EPTCS. Available at https://ntrs.nasa.gov/citations/. Note.

[13] A. Ut, H. Or & Co Author (1987): Title R2: A PhDthesis. Ph.D. thesis, School, Address, doi:10.4204/

EPTCS. Available at https://ntrs.nasa.gov/citations/20205011579. Note.

[14] A. Ut, H. Or & Co Author (1987): Title R3: A Masterthesis. Master’s thesis, School, Address, doi:10.4204/

EPTCS. arXiv:2042.12345. Note.

[15] A. Ut, H. Or & Co Author (1987): Title R4: A Manual, second edition. Organization, Address, doi:10.

4204/EPTCS. arXiv:2042.12345. Note.

[16] A. Ut, H. Or & Co Author (1987): Title U: A Booklet. Howpublished, Address, doi:10.4204/EPTCS. Avail-

able at https://ntrs.nasa.gov/citations/20205011579. Note.

[17] A. Ut, H. Or & Co Author (1987): Title U: A Misc. Howpublished, doi:10.4204/EPTCS. Available at

https://ntrs.nasa.gov/citations/20205011579. Note.

[18] A. Ut, H. Or & Co Author (1987): Title U: An Unpublished, doi:10.4204/EPTCS. arXiv:2042.12345. Note.